Table 1: L R HRS detector long checklist part 1. Completed by . Date:

Checked Needs attention
↓ ↓
1  The air cooling in the detector hut is operational
2  Each of the cooling fans (9 of them) are working in the electronics relay racks
3  Each of the cooling fans in the NIM crates (3 per crate) are working
4  The cooling fan is working on the VME crate
5  The cooling fan is working on the Trigger Supervisor Crate
6  The cooling fan is working in the VME slow control VME crate
7  Each of the cooling fans (3 of them) are working in the FastBus relay rack
8  The cooling fans are working on the VDC level translator panel
9  Power to the front-end discs is turned ON for the S2m and cooling fans are operational
10  Fresh air flow is turned on for PMTs on S0, S1 and S2m detectors, P ≈ 0.5 to 1 inches
11  All high voltage and signal cables in the detector hut are secured
12  The control of the MLU (VME) unit for the trigger is operational (via GUI)
13  The control of the HV crate for the detector is operational (via GUI)
14  The internet connection to each of the three FastBus CPUs is operational (via GUI)
15  The internet connection to the VME CPU is operational (via GUI)
16  The internet connection to the Trigger Supervisor CPU is operational (via GUI)
17  The VDC gas flows are turned on at a level of 30 in the flowmeters
18  The flow is seen from the exhaust bubblers/flowmeters
19  The power supply is turned on for the VDC amplifier cards
20  The threshold control for the VDC is turned on (+3V, 0.1A)
21  The level translators for the VDC are turned on (+5V and -5V)
22  The VDC HV power is set at 3.5 kV and the leakage current is below 0.01 μA (HV GUI)
23  The VDC leakage current trip level set to below 100 μA in HV GUI

1 Table 2: L R HRS detector long checklist part 2. Completed by . Date:

Checked Needs attention
↓ ↓
24  The Straw chamber gas flow is turned on at a level of 45 in the flowmeters
25  The pressure on the gauge is about 0.2 inches
26  The power supply is turned on for the Straws amplifier cards
27  The threshold control for the Straws is turned on
28  The level translators for the Straws are turned on
29  The Straw HV power is set at 1.6 kV and the leakage current is below 1 μA on each of the six lines (HV GUI)
30  The Straw chamber leakage current trip level set to below 200 μA in HV GUI
31  The S2m HV power is set at 1.6 kV settings and the current at each of the 32 channels is < 1 mA (HV GUI)
32  The S2m scaler rates are below 2 kHz in each of the 32 channels
33  The S0 HV power is set at 1.2 kV settings and the current at each of the 32 channels is < 1 mA (HV GUI)
34  The S0 scaler rates are below 2 kHz
35  The Gas Cherenkov, GC, CO2 flow is turned on at the level of 50 at the flowmeter
36  The GC HV power is set at 1.6 kV settings and the current at each of the 10 channels is < 1 mA (HV GUI)
37  The GC scaler rates are below 10 kHz in each of the 10 channels
38  The Shower HV power is set at 0.8 kV settings and the current at each of the channels is < 1 mA (HV GUI)
Date  Check performed by  Check approved (B. Wojtsekhowski)

2